OpenStudy (anonymous):

4x + y= -9(2,9) write an equation for the line that is parallel to the given line and that contains the given point?

OpenStudy (anonymous):

A parallel line has the same slope but different y intercept than the 1st equation. So, y = -4x + 9 --- 1st equation rearranged 9 = -4(2) + z ---- 2ns equation where z = y intercept solve for z z = 17 Therefore, the new equation is y = -4x + 17 or 17 = 4x + y

OpenStudy (anonymous):

Ok, well \[y=-9-4x\] so the slope is -4 for that line. So the parallel line must also have the slope -4. Now we have for the second line: \[y=-4x+b\] So all we need to do now is solve for b. Luckily we have a point given to us so \[9=-4(2)+b\] so b = 17

OpenStudy (anonymous):

-5x + 3y = 6;(-3,-8)

OpenStudy (anonymous):

i also need help with tish you if you dont mind!

OpenStudy (anonymous):

its good explanation

OpenStudy (anonymous):

but also a good question